Find the value of absolute zero on the Fahrenheit scale
1
Expert's answer
2019-07-29T08:18:39-0400

Absolute zero is 0 K.

"T [K] = 273 + 5/9 \cdot (T [\degree F] - 32)"

"T [\degree F] = 9/5 \cdot (T [K] - 273) + 32"

"T [\degree F] = 9/5 \cdot (0 - 273) + 32 = -459.4 \; \degree F"


Answer: "-459.4 \; \degree F".